# thanks: corresponding authorthanks: corresponding author

### v3

@@ 3 Discussion
Our numerical simulations demonstrated that our post-selection strategies achieve signifcant reductions in the logical error rate $p_{\mathrm{log}}$ (for accepted runs) by several orders of magnitude at moderate abort rates $p_{\mathrm{abort}}$. This performance gain is observed consistently across diverse code families: surface codes, bivariate bicycle codes, and hypergraph product codes (Figs. 24). For instance, by applying our strategy to the $[[144,12,12]]$ bivariate bicycle (BB) code, the logical error rate can be reduced by about three orders of magnitude with the abort rate of only around 1% (19%) at the physical error rate ($p$) of 0.1% (0.3%) under the uniform circuit-level noise model.
@@
Furthermore, real-time decoding is a key requirement for practical quantum computing involving non-Clifford gates, which should be sufficiently fast to avoid the backlog problem [1]. We extended our approach to such real-time scenarios through sliding-window decoding, featuring mid-circuit abort decisions that allow further computational cost reduction. The real-time strategy achieves performance comparable to (or even superior to) global post-selection in terms of the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$ for achieving a target logical error rate $p_{\mathrm{log}}$ (Fig. 5). Notably, the real-time strategy maintains favorable scaling with the number of rounds $T$, exhibiting nearly constant per-round logical error and abort rates (Fig. 6). In contrast, the global strategy does not preserve this consistency, making the real-time approach especially advantageous in practical regimes with large $T$.
@@
  • Further performance improvements: While cluster-based metrics offer practical advantages, the logical gap method still outperforms them for surface codes. Developing better metrics that approach logical gap performance while retaining computational efficiency and generalizability remains an open challenge.
@@
  • Practical applications: Evaluating the effectiveness of these strategies for specific quantum computing tasks, such as magic state preparation, would verify their practical utility.
@@ 4.1 Sliding-window decoding
We now describe the algorithm more formally. Consider a check matrix $H\in\{0,1\}^{r\times N}$ for $r$ detectors and $N$ error mechanisms. Let $\mathbf{s}\in\{0,1\}^{r}$ denote the syndrome vector and $\mathbf{p}\in[0,1]^{N}$ the prior probability vector for error mechanisms. We denote the sets of detector and error mechanism indices as $\mathcal{D}\coloneqq\{1,\cdots,r\}$ and $\mathcal{E}\coloneqq\{1,\cdots,N\}$, respectively. Each detector $i\in\mathcal{D}$ is assigned a time coordinate $t_{i}\in\{0,1,2,\cdots\}$. The algorithm maintains a global correction vector $\hat{\mathbf{e}}$, initially set to $\hat{\mathbf{e}}\leftarrow\mathbf{0}=(0,\cdots,0)\in\{0,1\}^{N}$.
@@
Next, the global state is updated by incorporating the committed corrections. To map the window-local solution back to the global problem, we define a commit vector $\hat{\mathbf{e}}_{w}^{\mathrm{commit}}\in\{0,1\}^{N}$ as
@@ Acknowledgements
S.H.L. conceived the idea and all authors contributed to developing it. S.H.L. conducted simulations and all authors interpreted the results together. S.H.L. mainly wrote the manuscript and L.E. made the first draft of Figure 1. All authors reviewed the manuscript.

### v4

@@ 3 Discussion
Our numerical simulations demonstrated that our post-selection strategies achieve significant reductions in the logical error rate $p_{\mathrm{log}}$ (for accepted runs) by several orders of magnitude at moderate abort rates $p_{\mathrm{abort}}$. This performance gain is observed consistently across diverse code families: surface codes, bivariate bicycle codes, and hypergraph product codes (Figs. 35). For instance, by applying our strategy to the $[[144,12,12]]$ bivariate bicycle (BB) code, the logical error rate can be reduced by about three orders of magnitude with the abort rate of only around 1% (19%) at the physical error rate ($p$) of 0.1% (0.3%) under the uniform circuit-level noise model. Compared to strategies based on other confidence metrics, the cluster-based approach significantly outperforms baseline metrics such as the correction weight and detector density across all code families. While the logical gap method with MWPM remains superior for surface codes, cluster-based metrics outperform the logical gap from BP+LSD or its proxy calculated by partial comparative decoding with a subset of logical classes.
@@
Furthermore, real-time decoding is essential for practical quantum computing with non-Clifford gates, which require Clifford corrections conditioned on decoded measurement outcomes [65]. Such real-time decoding must be sufficiently fast to avoid the exponential backlog problem. We extended our approach to this setting through sliding-window decoding, featuring mid-circuit abort decisions that terminate low-confidence trials early and save computational resources. The real-time strategy achieves performance comparable to, or even superior to, global post-selection in terms of the average time cost per accepted shot $\bar{T}_{\mathrm{accepted}}$ for achieving a target logical error rate $p_{\mathrm{log}}$ (Fig. 6). Notably, the real-time strategy exhibits favorable scaling with the number of rounds $T$, maintaining nearly constant per-round logical error and abort rates (Fig. 7). In contrast, the global strategy does not preserve this consistency; thus, the real-time approach offers more predictable and stable behavior for any $T$.
One may wonder whether real-time post-selection is truly necessary, given that for offline tasks (the main target of post-selection) backlog might not be the primary concern. We believe it remains valuable, for several reasons: (1) mid-circuit aborts can reduce expected resource costs by terminating low-confidence trials early, regardless of whether backlog is an issue; (2) even offline tasks may require real-time decoding when they involve logical non-Clifford gates that necessitate intermediate Clifford corrections (e.g., multi-round magic state distillation, where logical non-Clifford gates are implemented by consuming magic states distilled from earlier rounds); (3) decoder-confidence-based post-selection may also be useful for early fault-tolerant algorithms with moderate circuit depths, as suggested by recent studies on fault-tolerant error mitigation using decoder soft outputs [71, 27]; and (4) the favorable scaling with circuit depth demonstrated above can also benefit offline tasks with substantial depths, such as multi-round distillation.
@@
  • Further performance improvements: While cluster-based metrics offer practical advantages, the logical gap method still outperforms them for surface codes. Developing better metrics that approach logical gap performance while retaining computational efficiency and generalizability remains an open challenge. One promising direction could be incorporating geometric information about clusters beyond their sizes, such as inter-cluster distances. However, inter-cluster distance alone would be insufficient: when only a single cluster exists, this distance is undefined, yet confidence can still vary significantly depending on the cluster’s size. One might also consider distances between clusters and code boundaries, but this raises additional challenges for codes with periodic boundary conditions or general QLDPC codes where the notion of a boundary may not be rigorously defined. Designing optimal confidence metrics that appropriately incorporate such distance information remains a nontrivial open problem.
  • Gap proxy methods for multi-qubit codes: Our gap proxy analysis (Supplementary Note 1) shows that various heuristics for selecting a subset of logical classes to decode do not outperform naive random sampling, and the resulting metric fails to outperform the correction weight method despite significant computational overhead. This is because post-selection performance depends on the discrimination power between successful and failed samples, not approximation accuracy; distribution-based heuristics improve accuracy primarily for successful shots, thereby reducing discriminative power. Effective gap proxy methods may require fundamentally different approaches.
@@
  • Practical applications: Evaluating the effectiveness of these strategies for specific quantum computing tasks, such as magic state preparation, would verify their practical utility. It would also be interesting to test whether these metrics are useful as decoder-confidence signals for logical-level error mitigation, analogous to recent studies using decoder soft information [71, 27, 1].
@@ 4 Methods
4.1 Formal definitions of the graph-theoretic structures on clustering-based decoders
In this subsection, we provide formal definitions of the graph-theoretic structures underlying clustering-based decoders, as introduced in the Results section.
For an $n$-dimensional vector $\mathbf{v}$, we denote its $i$-th entry by $v_{i}$ and, for an index set $I\subseteq\{1,\dots,n\}$, we write $\mathbf{v}[I]$ for the subvector consisting of $\{v_{i}:i\in I\}$, ordered by increasing indices. For an $n\times m$ matrix $M$, we denote its $(i,j)$-th entry by $M[i,j]$ and, for $I_{1}\subseteq\{1,\dots,n\}$ and $I_{2}\subseteq\{1,\dots,m\}$, we write $M[I_{1},I_{2}]$ for the submatrix with rows indexed by $I_{1}$ and columns indexed by $I_{2}$.
A detector error model (DEM) specifies a set of error mechanisms $\mathcal{E}=\{1,\ldots,N\}$, a set of detectors $\mathcal{D}=\{1,\ldots,r\}$, and the relationships between them. The check matrix $H\in\mathbb{Z}_{2}^{r\times N}$ is a binary matrix where $H[i,j]=1$ if and only if detector $i$ is flipped by error mechanism $j$. Each error mechanism $j$ has an associated prior probability $p_{j}\in[0,1]$.
Definition 3 (Tanner graph).
The Tanner graph $\mathcal{G}_{T}=(\mathcal{D}\cup\mathcal{E},E_{T})$ is a bipartite graph where the vertex set is partitioned into check nodes (corresponding to detectors $\mathcal{D}$) and fault nodes (corresponding to error mechanisms $\mathcal{E}$). An edge $(i,j)\in E_{T}$ exists between check node $i\in\mathcal{D}$ and fault node $j\in\mathcal{E}$ if and only if $H[i,j]=1$. The check matrix $H$ thus serves as the biadjacency matrix of the Tanner graph.
Given the Tanner graph structure, we can formally state the decoding problem.
Definition 4 (Decoding problem).
Given a check matrix $H\in\mathbb{Z}_{2}^{r\times N}$ and a syndrome $\mathbf{s}\in\mathbb{Z}_{2}^{r}$ (a binary vector indicating which detectors are violated), the decoding problem is to find a correction $\tilde{\mathbf{e}}\in\mathbb{Z}_{2}^{N}$ such that
$H\tilde{\mathbf{e}}=\mathbf{s}.$
In general, multiple solutions exist; a decoder selects one according to a specific strategy (e.g., minimizing the log-likelihood weight).
While the Tanner graph captures the relationship between error mechanisms and detectors, it is often useful to consider the connectivity structure among error mechanisms directly. This motivates the construction of the fault graph.
Definition 5 (Fault graph).
The fault graph $\mathcal{G}_{F}=(\mathcal{E},E_{F})$ is constructed by projecting the Tanner graph onto the fault nodes. Specifically, two fault nodes $j_{1},j_{2}\in\mathcal{E}$ (with $j_{1}\neq j_{2}$) are connected by an edge in $E_{F}$ if and only if there exists at least one detector $i\in\mathcal{D}$ such that $H[i,j_{1}]=H[i,j_{2}]=1$. In other words, two error mechanisms are adjacent in the fault graph if they share a common detector. The fault graph can equivalently be characterized through the Gram-like matrix $H^{\top}H$: an edge exists between $j_{1}$ and $j_{2}$ if and only if $(H^{\top}H)[j_{1},j_{2}]\geq 1$.
Clustering-based decoders operate by partitioning error mechanisms into groups that can be decoded independently.
Definition 6 (Cluster).
A cluster $C\subseteq\mathcal{E}$ is a set of error mechanisms that forms a connected subgraph in the fault graph $\mathcal{G}_{F}$. Equivalently, for any two error mechanisms $j_{1},j_{2}\in C$, there exists a path in $\mathcal{G}_{F}$ connecting $j_{1}$ to $j_{2}$ that passes only through vertices in $C$.
For a cluster to be useful for decoding, it must be “self-contained” in the sense that its internal structure suffices to determine a consistent correction. This requirement is formalized through the notion of validity.
Definition 7 (Valid cluster).
Let $C\subseteq\mathcal{E}$ be a cluster, and let $\mathcal{D}_{C}\coloneqq\{i\in\mathcal{D}:\exists j\in C\text{ such that }H[i,j]=1\}$ denote the set of detectors that involve at least one error mechanism in $C$. The cluster $C$ is valid if there exists a correction $\tilde{\mathbf{e}}\in\mathbb{Z}_{2}^{N}$ such that for every detector $i\in\mathcal{D}_{C}$,
$\bigoplus_{j\in C}H[i,j]\cdot\tilde{e}_{j}=s_{i},$
where $s_{i}\in\mathbb{Z}_{2}$ is the observed syndrome value for detector $i$, and $\oplus$” denotes addition modulo 2. In other words,
$\mathbf{s}[\mathcal{D}_{C}]\in\operatorname{IMAGE}(H[\mathcal{D}_{C},C]).$ (3)
Intuitively, a valid cluster contains sufficient error mechanisms to explain all syndrome bits that it touches. The validity condition ensures that the cluster can be decoded in isolation without affecting or being affected by error mechanisms outside the cluster.
We now formally define the class of decoders that leverage this cluster structure.
Definition 8 (Clustering-based decoder).
A clustering-based decoder is a decoder that solves the decoding problem (Definition 4) through the following procedure:
  1. 1.
    Cluster construction: Construct a collection of disjoint valid clusters $\{C_{1},C_{2},\ldots,C_{m}\}$ such that $\bigcup_{k=1}^{m}\mathcal{D}_{C_{k}}$ covers all violated detectors.
  2. 2.
    Local decoding: For each cluster $C_{k}$, solve the local decoding problem restricted to the submatrix $H[\mathcal{D}_{C_{k}},C_{k}]$ to obtain a local correction $\tilde{\mathbf{e}}_{k}$.
  3. 3.
    Correction aggregation: Combine the local corrections to form the global correction $\tilde{\mathbf{e}}$, where $\tilde{e}_{j}=\tilde{e}_{k,j}$ if $j\in C_{k}$ for some $k$, and $\tilde{e}_{j}=0$ otherwise.
The efficiency of clustering-based decoders stems from the fact that small, localized clusters require less computational effort to decode than solving the full decoding problem globally and can benefit from parallelization.
4.2 Belief propagation plus localized statistics decoding (BP+LSD)
We briefly describe the BP+LSD decoder [39] using the notation established in the preceding subsection. Given a check matrix $H\in\mathbb{Z}_{2}^{r\times N}$ and syndrome $\mathbf{s}\in\mathbb{Z}_{2}^{r}$, the decoder aims to find a correction $\tilde{\mathbf{e}}\in\mathbb{Z}_{2}^{N}$ satisfying $H\tilde{\mathbf{e}}=\mathbf{s}$. BP+LSD first attempts to solve this problem using belief propagation (BP), and if BP fails to converge, it applies localized statistics decoding (LSD) as a postprocessing step.
Belief propagation phase.
BP is executed on the Tanner graph $\mathcal{G}_{T}$ (Definition 3) for a fixed number of iterations. This produces a posterior LLR $\ell_{j}\in\mathbb{R}$ for each error mechanism $j\in\mathcal{E}$, where smaller $\ell_{j}$ indicates higher posterior probability that the error mechanism occurred. If BP converges to a valid solution $\tilde{\mathbf{e}}_{\mathrm{BP}}$ satisfying $H\tilde{\mathbf{e}}_{\mathrm{BP}}=\mathbf{s}$, this solution is returned. Otherwise, the posterior LLRs $\{\ell_{j}\}_{j\in\mathcal{E}}$ are passed to the LSD phase as reliability guidance.
LSD cluster initialization, growth, and merging.
LSD constructs valid clusters (Definition 7) by growing them from violated detectors. For each detector $i$ with $s_{i}=1$ (violated), an initially empty cluster $C$ is created with $\mathcal{D}_{C}=\{i\}$. Each cluster is then grown incrementally by adding exactly one error mechanism per step. To select which error mechanism to add, we define the boundary detectors $\beta(C):=\{i\in\mathcal{D}_{C}:\exists j\notin C\text{ with }H[i,j]=1\}$ and the candidate error mechanisms $\Lambda(C):=\{j\notin C:\exists i\in\beta(C)\text{ with }H[i,j]=1\}$. At each step, the error mechanism with the smallest posterior LLR among candidates is added:
$j^{\star}:=\arg\min_{j\in\Lambda(C)}\ell_{j},\qquad C\leftarrow C\cup\{j^{\star}\}.$ (4)
When $j^{\star}$ connects to boundary detectors of multiple clusters, these clusters are merged by taking the union of their error mechanism sets.
Validity check.
After each growth or merging step, the algorithm checks whether the updated cluster is valid, i.e., whether the local syndrome lies in the image of the local check matrix [Eq. (3)]. This is implemented via on-the-fly elimination: the decoder maintains an incremental elimination state (a row-echelon form over $\mathbb{Z}_{2}$) for $H[\mathcal{D}_{C},C]$ and applies the same stored row operations to $\mathbf{s}[\mathcal{D}_{C}]$, yielding a reduced augmented system $U\mid\mathbf{s}^{\prime}]$. When the cluster grows by adding a new error mechanism, only the newly added column is eliminated using the stored operations (cluster merging is handled similarly). The cluster is valid if and only if $U\mid\mathbf{s}^{\prime}]$ contains no row of the form $0\cdots 0\mid 1]$.
Local decoding and global assembly.
Growth and merging continue until all clusters become valid. For each valid cluster $C$, a local solution $\tilde{\mathbf{e}}_{C}$ satisfying $H[\mathcal{D}_{C},C]\tilde{\mathbf{e}}_{C}=\mathbf{s}[\mathcal{D}_{C}]$ is computed using the maintained factorization. The global correction $\tilde{\mathbf{e}}$ is assembled by placing each local solution on its corresponding coordinates and setting all other entries to zero.
@@ 4.3 Sliding-window decoding
We now describe the algorithm more formally, using the notation introduced previously: the check matrix $H\in\mathbb{Z}_{2}^{r\times N}$, the sets of detectors $\mathcal{D}$ and error mechanisms $\mathcal{E}$, and the syndrome vector $\mathbf{s}\in\mathbb{Z}_{2}^{r}$. Additionally, each detector $i\in\mathcal{D}$ is assigned a time coordinate $t_{i}\in\{0,1,2,\cdots\}$. The algorithm maintains a global correction vector $\hat{\mathbf{e}}$, initially set to $\hat{\mathbf{e}}\leftarrow\mathbf{0}=(0,\cdots,0)\in\mathbb{Z}_{2}^{N}$.
@@
Next, the global state is updated by incorporating the committed corrections. To map the window-local solution back to the global problem, we define a commit vector $\hat{\mathbf{e}}_{w}^{\mathrm{commit}}\in\mathbb{Z}_{2}^{N}$ as
@@ Acknowledgements
S.H.L. conceived the idea and all authors contributed to developing it. S.H.L. conducted simulations and all authors interpreted the results together. S.H.L. mainly wrote the manuscript, and L.E. and S.B. contributed helpful feedback and revisions. All authors reviewed the manuscript.